Find the global maximum of 2 = f(x,y) = 3y - xy over the region bounded by y=x², y = 0, and x = 4.
Answer #1

Given solutions stepli flny) = 34.xy 9=22 - 0 gcn, y) = y.x² Cet By using Lagrange multipliers we get of a xog <-4, 3-x 72 xcat (0,0) : f(0,0) = 0 at Ca, u) : f (2,4)= 3(4)- 2(4) 71-8 =) ( f (8,624) of (4160) : f(4,16) = 3016) - 4 (16) - 48-60 - (f